# Noncommutative weak type $(1,1)$ estimates for Calderón-Zygmund operators with $L_{2}$-mean Hörmander conditions

### v1

@@ 2. $(ii)$
$\int_{\operatorname{\mathbb{R}}^{d}}\Big\|\int_{B_{r_{n-j}}}\int_{Q}\big(K_{j}(x,y+z)-K_{j}(x,y)\big)b_{\text{off},n-j}^{Q,i}(y,z){\rm d}y{\rm d}z\Big\|_{L_{1}(\operatorname{\mathcal{M}})}{\rm d}x,$
for $i=1,2,3,4$, with
$b_{\text{off},n-j}^{Q,1}(y,z)=p_{Q}(f\chi_{Q})(y)h_{r_{n-j}}(z)q_{Q},\quad b_{\text{off},n-j}^{Q,2}(y,z)=p_{Q}(f_{Q}\chi_{Q})(y)h_{r_{n-j}}(z)q_{Q},$
$b_{\text{off},n-j}^{Q,3}(y,z)=q_{Q}(f\chi_{Q})(y)h_{r_{n-j}}(z)p_{Q},\quad b_{\text{off},n-j}^{Q,4}(y,z)=q_{Q}(f_{Q}\chi_{Q})(y)h_{r_{n-j}}(z)p_{Q}.$
Again, in the following we only estimate $G_{1}$ in detail; similar arguments apply to $G_{2}$$G_{4}$ by using $\phi(f_{n}p_{n})=\phi(fp_{n})$.
Consider $G_{1}$. It follows from Lemma A.3 (in Appendix A), (2.1), Lemma 4.4 $(ii)$, and the $L_{2}$-mean Hörmander condition (1.7) that
$G_{1}$ $\lesssim\sum_{n\geq 2}\sum_{j\leq n-1}\sum_{Q\in\operatorname{\mathcal{Q}}_{n-j}}(|Q|\lambda)^{\frac{1}{2}}$
$\ \ \ \ \frac{1}{|B_{r_{n-j}}|}\int_{\operatorname{\mathbb{R}}^{d}}\Big\|\int_{Q}\Big(\int_{B_{r_{n-j}}}\left|K_{i}(x,y+z)-K_{i}(x,y)\right|{\rm d}z\Big)^{2}p_{Q}f(y)p_{Q}{\rm d}y\Big\|_{L_{\frac{1}{2}}(\operatorname{\mathcal{M}})}^{\frac{1}{2}}{\rm d}x$
$\lesssim\sum_{n\geq 2}\sum_{j\leq n-1}\sum_{Q\in\operatorname{\mathcal{Q}}_{n-j}}\big(2H_{2}(n)+2^{-n}\big)\big(\tau(p_{Q})\phi(fp_{Q}\chi_{Q})|Q|\lambda\big)^{\frac{1}{2}}$
$\lesssim\sum_{m\geq 1}\sum_{Q\in\operatorname{\mathcal{Q}}_{m}}\big(\tau(p_{Q})\phi(fp_{Q}\chi_{Q})|Q|\lambda\big)^{\frac{1}{2}}.$
Then, the remaining part of the proof follows the same method as used for $F_{1}$. This completes the estimate for $G_{1}$ and the proof of Lemma 4.2. ∎
@@ Appendix A Noncommutative $L_{p}$ spaces
This part consists of three subsections: noncommutative $L_{p}$ spaces, noncommutative maximal norms, and noncommutative square functions. All these content are standard in noncommutative harmonic analysis, we refer to see [Cuculescu1971Martingales, Junge2002Doob, Parcet2009Pseudo-localization, Pisier1998noncommutative, PisierQuanhua2003] for more details.
@@ A.1. Noncommutative $L_{p}$ spaces
The noncommutative $L_{p}$ space associated with $(\operatorname{\mathcal{M}},\tau)$, denoted by $L_{p}(\operatorname{\mathcal{M}},\tau)$ or simply $L_{p}(\operatorname{\mathcal{M}})$, is defined as the completion of the space $(\operatorname{\mathcal{S}},\left\|\cdot\right\|_{L_{p}(\operatorname{\mathcal{M}})})$. For convenience, we set $L_{\infty}(\operatorname{\mathcal{M}})=\operatorname{\mathcal{M}}$ equipped with the operator norm $\left\|\cdot\right\|_{\operatorname{\mathcal{M}}}$. We refer the reader to [PisierQuanhua2003] for more information regarding noncommutative $L_{p}$ spaces.
@@ A.2. Noncommutative maximal norms
The generalization of the maximal function to noncommutative spaces faces a challenge: operators in von Neumann algebras are not directly comparable. Remarkably, Pisier [Pisier1998noncommutative] and Junge [Junge2002Doob] overcame this difficulty by directly defining the noncommutative maximal norm.
@@ A.3. Noncommutative square functions
Then, the column space $L_{p}(\operatorname{\mathcal{M}};L_{2}^{c}(\Omega))$ is defined as the space of all such functions $f$ for which $\left\|f\right\|_{L_{p}(\operatorname{\mathcal{M}};L_{2}^{c}(\Omega))}<\infty$. The column space is the foundation for the following Hölder inequality, which will be useful later. We omit the symbol $\mu$ when no ambiguity arises.
@@
We refer the reader to [PisierQuanhua2003] for more information about Hilbert valued operator spaces.

### v2

@@ 2. $(ii)$
$\int_{\operatorname{\mathbb{R}}^{d}}\Big\|\int_{Q}\int_{|z|\leq\frac{\sqrt{d}}{2}\ell(Q)}\big(K_{j}(x,y+z)-K_{j}(x,y)\big)b_{\text{off},n-j}^{Q,i}(y,z){\rm d}z{\rm d}y\Big\|_{L_{1}(\operatorname{\mathcal{M}})}{\rm d}x,$
for $i=1,2,3,4$, where
$b_{\text{off},n-j}^{Q,1}(y,z)=p_{Q}(f\chi_{Q})(y)h_{\frac{\sqrt{d}}{2}\ell(Q)}(z)q_{Q},\quad b_{\text{off},n-j}^{Q,2}(y,z)=p_{Q}(f_{Q}\chi_{Q})(y)h_{\frac{\sqrt{d}}{2}\ell(Q)}(z)q_{Q},$
$b_{\text{off},n-j}^{Q,3}(y,z)=q_{Q}(f\chi_{Q})(y)h_{\frac{\sqrt{d}}{2}\ell(Q)}(z)p_{Q},\quad b_{\text{off},n-j}^{Q,4}(y,z)=q_{Q}(f_{Q}\chi_{Q})(y)h_{\frac{\sqrt{d}}{2}\ell(Q)}(z)p_{Q}.$
We estimate only $G_{1}$; the other terms are similar by using $\phi(f_{n}p_{n})=\phi(fp_{n})$.
Fix a cube $Q\in\operatorname{\mathcal{Q}}_{n-j}$. By the supports of $h_{\frac{\sqrt{d}}{2}\ell(Q)}$ and $\chi_{Q}$, the noncommutative $L_{1}$-norm inside $G_{1}$ equals
$\Big\|\int_{Q}\mathcal{J}_{j}(x,y)p_{Q}f(y)q_{Q}{\rm d}y\Big\|_{L_{1}(\operatorname{\mathcal{M}})},$
where
$\mathcal{J}_{j}(x,y):=\int_{|z|\leq\frac{\sqrt{d}}{2}\ell(Q)}\big(K_{j}(x,y+z)-K_{j}(x,y)\big)h_{\frac{\sqrt{d}}{2}\ell(Q)}(z){\rm d}z.$
We also write $\mathcal{J}_{j}(x,y)=\mathcal{J}^{+}_{j}(x,y)-\mathcal{J}^{-}_{j}(x,y)$ and consider only $\mathcal{J}^{+}_{j}(x,y)$, i.e.
(4.13) $\Big\|\int_{Q}\mathcal{J}_{j}^{+}(x,y)p_{Q}f(y)q_{Q}{\rm d}y\Big\|_{L_{1}(\operatorname{\mathcal{M}})},$
Immediately, when $y\in Q$, it follows from the definition of $h_{\frac{\sqrt{d}}{2}\ell(Q)}$ that
$|\mathcal{J}^{+}_{j}(x,y)|\lesssim\frac{1}{|Q|}\int_{|z|\leq\frac{\sqrt{d}}{2}\ell(Q)}|K_{j}(x,y+z)-K_{j}(x,y)|{\rm d}z\leq\sup_{y\in Q}\mathcal{K}_{j,Q}(x,y).$
Now, following the same reasoning as for $F_{1}$ in Lemma 4.2 $(i)$, we bound (4.13) by
$\Big\|\Big(\int_{Q}\mathcal{J}_{j}^{+}(x,y)p_{Q}f(y)p_{Q}{\rm d}y\Big)^{\frac{1}{2}}\Big\|_{L_{2}(\operatorname{\mathcal{M}})}\Big\|p_{Q}m^{\prime}(x)\Big(\int_{Q}\mathcal{J}_{j}^{+}(x,y)q_{Q}f(y)q_{Q}{\rm d}y\Big)^{\frac{1}{2}}\Big\|_{L_{2}(\operatorname{\mathcal{M}})}$
$\lesssim\Big(\sup_{y\in Q}\mathcal{K}_{j,Q}(x,y)\Big)\tau\Big(\int_{Q}p_{Q}f(y)p_{Q}{\rm d}y\Big)^{\frac{1}{2}}\tau\Big(p_{Q}m^{\prime}(x)\big(\int_{Q}q_{Q}f(u)q_{Q}{\rm d}u\big)(m^{\prime}(x))^{*}p_{Q}\Big)^{\frac{1}{2}},$
where $m^{\prime}(x)$ is the contraction provided by Lemma A.4 in Appendix A.3. Finally, the remaining part of the proof follows the same method as used for $F_{1}$. Hence, we complete the estimate for $G_{1}$ and the proof of Lemma 4.2.
@@ Appendix A Noncommutative $L_{p}$ spaces
This part consists of three subsections: noncommutative $L_{p}$ spaces, noncommutative maximal norms, and noncommutative square functions. All these content are standard in noncommutative harmonic analysis, we refer to see [5, 9, 17, 18, 19] for more details.
@@ A.1. Noncommutative $L_{p}$ spaces
The noncommutative $L_{p}$ space associated with $(\operatorname{\mathcal{M}},\tau)$, denoted by $L_{p}(\operatorname{\mathcal{M}},\tau)$ or simply $L_{p}(\operatorname{\mathcal{M}})$, is defined as the completion of the space $(\operatorname{\mathcal{S}},\left\|\cdot\right\|_{L_{p}(\operatorname{\mathcal{M}})})$. For convenience, we set $L_{\infty}(\operatorname{\mathcal{M}})=\operatorname{\mathcal{M}}$ equipped with the operator norm $\left\|\cdot\right\|_{\operatorname{\mathcal{M}}}$. We refer the reader to [19] for more information regarding noncommutative $L_{p}$ spaces.
@@ A.2. Noncommutative maximal norms
The generalization of the maximal function to noncommutative spaces faces a challenge: operators in von Neumann algebras are not directly comparable. Remarkably, Pisier [18] and Junge [9] overcame this difficulty by directly defining the noncommutative maximal norm.
@@ A.3. Noncommutative square functions
Then, the column space $L_{p}(\operatorname{\mathcal{M}};L_{2}^{c}(\Omega))$ is defined as the space of all such functions $f$ for which $\left\|f\right\|_{L_{p}(\operatorname{\mathcal{M}};L_{2}^{c}(\Omega))}<\infty$. When the measure is clear, we write simply ${\rm d}\omega$ instead of ${\rm d}\mu(\omega)$.
First, we introduce a Hölder inequality for the square functions.
@@
The next result was communicated to the authors of [7] by Cadilhac during their work on the maximal Calderón-Zygmund operators. We include a proof for completeness.
Lemma A.4.
Let $0<p,q\leq\infty$. For any $f\in L_{p}(\operatorname{\mathcal{M}};L_{2}^{c}(\Sigma))$ and $g\in L_{q}(\operatorname{\mathcal{M}};L_{2}^{c}(\Sigma))$, there exists a contraction operator $m$ such that
$\int_{\Sigma}f^{*}(\omega)g(\omega){\rm d}\omega=\Big(\int_{\Sigma}f^{*}(\omega)f(\omega){\rm d}\omega\Big)^{\frac{1}{2}}m\Big(\int_{\Sigma}g^{*}(\omega)g(\omega){\rm d}\omega\Big)^{\frac{1}{2}}.$
Proof.
Fix $\epsilon>0$ and set
$F_{\epsilon}:=\int_{\Sigma}f^{*}(\omega)f(\omega){\rm d}\omega+\epsilon{\textbf{1}}_{\operatorname{\mathcal{M}}}\quad\text{and}\quad G_{\epsilon}:=\int_{\Sigma}g^{*}(\omega)g(\omega){\rm d}\omega+\epsilon{\textbf{1}}_{\operatorname{\mathcal{M}}}.$
Then, we define $m_{\epsilon}$ by
$m_{\epsilon}:=F_{\epsilon}^{-\frac{1}{2}}\Big(\int_{\Sigma}f^{*}(\omega)g(\omega){\rm d}\omega\Big)G_{\epsilon}^{-\frac{1}{2}},$
so that
(A.1) $\int_{\Sigma}f^{*}(\omega)g(\omega){\rm d}\omega=F_{\epsilon}^{\frac{1}{2}}m_{\epsilon}G_{\epsilon}^{\frac{1}{2}}.$
[…]
$m_{\epsilon}=\int_{\Sigma}\big(f(\omega)F_{\epsilon}^{-\frac{1}{2}}\big)^{*}\big(g(\omega)G_{\epsilon}^{-\frac{1}{2}}\big){\rm d}\omega.$
Applying Lemma A.3, we get
$\|m_{\epsilon}\|_{L_{\infty}(\operatorname{\mathcal{M}})}$ $\leq\Big\|F_{\epsilon}^{-\frac{1}{2}}\Big(\int_{\Sigma}|f(\omega)|^{2}{\rm d}\omega\Big)^{\frac{1}{2}}F_{\epsilon}^{-\frac{1}{2}}\Big\|_{L_{\infty}(\operatorname{\mathcal{M}})}\Big\|G_{\epsilon}^{-\frac{1}{2}}\Big(\int_{\Sigma}|g(\omega)|^{2}{\rm d}\omega\Big)^{\frac{1}{2}}G_{\epsilon}^{-\frac{1}{2}}\Big\|_{L_{\infty}(\operatorname{\mathcal{M}})}$
$=\big\|F_{\epsilon}^{-\frac{1}{2}}F_{0}F_{\epsilon}^{-\frac{1}{2}}\big\|_{L_{\infty}(\operatorname{\mathcal{M}})}\big\|G_{\epsilon}^{-\frac{1}{2}}G_{0}G_{\epsilon}^{-\frac{1}{2}}\big\|_{L_{\infty}(\operatorname{\mathcal{M}})}$
$\leq 1,$
[…]
$\int_{\Sigma}f^{*}(\omega)g(\omega){\rm d}\omega=F_{0}^{\frac{1}{2}}mG_{0}^{\frac{1}{2}},$
which is the desired factorization. ∎
We refer the reader to [19] for more information about Hilbert valued operator spaces.
Acknowledgements
We thank L. Cadilhac for his comment about the $L_{q}$-mean Dini condition, which helped us improve this paper.